OpenStudy (anonymous):

Which fraction in simplest form is equivalent to zero point four? Answer four-sixteenths twelve-thirds eighty thirty-seconds twenty-four sixtieths

OpenStudy (alexwee123):

0.4=4/10=2/5

OpenStudy (alexwee123):

only one place value after the decimal

OpenStudy (anonymous):

seriously... did you even read the question? i am sorry, that is not correct, I mean it is right, just not in this context

OpenStudy (alexwee123):

yes i gave you a fraction of 0.4 so now compare it to the choices that you were given

OpenStudy (anonymous):

oh god...

OpenStudy (alexwee123):

?

OpenStudy (alexwee123):